# Alchemical harmonic approximation based potential for all iso-electronic diatomics: Foundational baseline for $\Delta$-machine learning

### v2

@@ Abstract
We introduce the alchemical harmonic approximation (AHA) of the absolute electronic energy for charge-neutral iso-electronic diatomics at fixed interatomic distance $d_{0}$. To account for variations in distance, we combine AHA with this Ansatz for the electronic binding potential, $E(d)=(E_{u}-E_{s})\left(\frac{E_{c}-E_{s}}{E_{u}-E_{s}}\right)^{\sqrt{d/d_{0}}}+E_{s}$, where $E_{u},E_{c},E_{s}$ correspond to the energies of united atom, calibration at $d_{0}$, and sum of infinitely separated atoms, respectively. Our model covers the entire two-dimensional electronic potential energy surface spanned by distance and difference in nuclear charge from which only one single point (with elements of nuclear charge $Z_{1},Z_{2}$ and distance $d_{0}$) is drawn to calibrate $E_{c}$. Using reference data from pbe0/cc-pVDZ, we present numerical evidence for the electronic ground-state of all neutral diatomics with 8, 10, 12, 14 electrons. We assess the validity of our model by comparison to legacy interatomic potentials (Harmonic oscillator, Lennard-Jones, and Morse) within the most relevant range of binding (0.7-2.5 Å), and find comparable accuracy if restricted to single diatomics, and significantly better predictive power when extrapolating to the entire iso-electronic series. We also investigated $\Delta$-learning of the electronic absolute energy using our model as baseline. This baseline model results in a systematic improvement, effectively reducing training data needs for reaching chemical accuracy by up to an order of magnitude from $\sim$1000 to $\sim$100. By contrast, using AHA+Morse as a baseline hardly leads to any improvement, and sometimes even deteriorates the predictive power. Inferring the energy of unseen CO converges to a prediction error of $\sim$0.1 Ha in direct learning, and $\sim$0.04 Ha with our baseline.
@@ II.2 The alchemical harmonic approximation (AHA)
From these three constraints, we find for the energy of the AHA:
@@ II.3 The Alchemical Quartic Approximation (AQuA)
Knowledge of $F_{c}$ gives a fourth constraint on the functional form of $E(\lambda)$. For reasons of symmetry, the next higher order in $E(\lambda)$ must be quartic, so $F_{c}$ together with Eqs. 2 and 3 should allow an implementation of the Alchemical Quartic Approximation (AQuA). However, solving the linear system of equations for a quartic polynomial necessitates to compute $\partial E_{s}/\partial\lambda$ (cf. Sec. II.4), introducing a numerical source of error. This error scales with $\lambda^{3}$ by construction, in addition to errors in the electron density (which in turn is built from basis sets optimized on energies of one element at question), rendering predictions from $F_{c}$ as in Eq. 5 prohibitive.
To include additional constraints into general functional forms of $E(\lambda)$, its zeros need to be found (cf. Eqs. 2, 3). The quadratic (AHA) and quartic models (AQuA) are readily available in this regard, but more complicated methods, e.g. the diatomic energy formula from Ref. michael_hammett_JACS, with powers of $\lambda^{7/3}$, do not allow such analytical treatment.
However, even when restricted to the AHA, we could have picked a different triplet of points or derivatives to determine the parabola instead of Eqs. 2, 3, e.g.
@@
However, both sets of three points rely on the alchemical force $F_{0}(d)$, i.e. any implementation necessitates the calculation of Eq. 18’s derivative w.r.t. $\lambda$ (see below) in addition to $F_{c}$. For numerical stability, we consider only Eq. 4 because its distance dependence can be expressed without derivatives in $\lambda$.
@@ II.4 Distance-dependence
To model the distance-dependence at some fixed $\lambda_{0}$, we have identified by trial and error the interatomic potential,
@@
The energies of the neutral atoms $Z_{1},Z_{2},Z_{u}$ are neither $d$-, nor $\lambda$-dependent, and can be precomputed.
@@
Implicit in this equation is the physical constraint $E_{u}<E_{c}<E_{s}$, i.e. the calibration must never be the united atom which was already a condition for the AHA/AQuA) but also never the infinitely (or very far) separated system! With Eq. 18, only one calibration point $E_{c}$ at $\lambda_{0},d_{0}$ is needed to determine the entire energy behavior $E(\lambda,d)$ of iso-electronic diatomics.
@@ III.1 Comparing potentials
Note that conventional potentials, such as Morse, Lennard-Jones or the harmonic potential (harmonic in $d$) all model distance-dependent nuclear repulsion and electronic attraction togther—despite the fact that the Coulombic part could have been easily subtracted in order to better focus on the contributions from quantum mechanics (Fig. 3). Especially in the short range, difficulties arise as Coulombic and empirical repulsion do not cancel exactly. The advantage of Eq. 18 is its correct and systematic behavior for $d=0$ and $d\rightarrow\infty$, and its dependence on only one calibration point, almost regardless of choice of $d$ and $\lambda$.
Computing parameters via fitting is not necessary for reasonable MAEs of our potential, but naturally, one can consider the quantity $E_{s}$ as fitting parameter to improve Eq. 18’s accuracy. In Fig. 4, we present a comparison of our potential with established potentials like the harmonic oscillator (HO), Lennard-Jones (LJ) and the Morse potential, all within the AHA. As calibration point, we choose BF ($\lambda_{0}=2$) and $d_{0}=1.2$. In the fitted potentials (HO, LJ, Morse, our(fit)), there is a clear improvement towards the calibration calculation at $\lambda=2$, but for larger deviations from the symmetric diatomic, the approximation of the AHA worsens (seen in the MAE of ours and ours(fit) which only predict electronic energies). The extreme changes of the other series (HO, LJ, Morse) in MAE stem from a similar source: an improvement of the MAE for unsymmetric charges $Z_{1},Z_{2}$ can be achieved in the fitting routine until the AHA approximation worsens to such a degree that all three potentials can no longer adequately describe the energy shape. Similar behavior can be found throughout the different calibrations (N2, CO, BeNe, LiNa, HeMg, HAl) and the series’ of 8 (Be2, LiB, HeC, HN), 10 (B2, BeC, LiN, HeO, HF) and 12 (C2, BN, BeO, LiF, HeNe, HNa) electrons.
@@ III.2 Machine Learning with AHA as baseline
As a calibration calculation, we again pick BF. Then, the difference between ours and the Morse potential as a baseline model for the prediction of diatomic energies and the reference data at DFT-level of theory (pbe0/cc-pVDZ) can be learned with KRR. For this, we generate 7 times 1024 points (for the 7 different $\lambda$ and 1024 points for $d\in[0.7,2.5]$   Å), then choose different global and local representations for the diatomics (see Sec. Computational Details). The kernel is Laplacian with the Manhattan norm. To determine hyperparameters, we employ 8-fold cross-validation for different training set sizes $N_{\text{train}}=8,16,32,64\dots,4096$. The size of the test set is $N_{\text{test}}=3072$. We compare these two $\Delta$-learning approaches with direct learning of the CM(n) representation (Fig. 5). Note the different sizes of representations: while CM(n) contains only four numbers per diatomic, cMBDF consists of 40 (times 2 atoms), FCHL19 of 3952 (times 2 atoms) and finally SOAP with 44296 (times 2 atoms)! Especially SOAP scales costly as every element species of the 14-electron diatomics is included, i.e. elements H to Si.
@@ III.3 Out-of-sample prediction of diatomics
We want to determine the MAE of the KRR model when systematically removing one specific diatomic, e.g. CO, from the training set to use as test set. Again, consider the AHA+ours and AHA+Morse baseline models with BF as calibration calculation. Instead of randomly selecting $N_{\text{train}}$ samples, restrict the data in the training set to $\lambda\neq 1$ and in the test set to $\lambda=1$. Since local representations do not work anymore (cf. Eq. 24), we exclusively consider the performance between the CM(n) representations in direct and $\Delta$-learning (inset of Fig. 5). Clearly, all three models approach an average energy prediction early but no accuracy is gained upon further increase of the training set size. However, the baseline model of AHA+ours gives a significant edge over the direct learning or $\Delta$-learning with AHA+Morse. Both the stagnation, as well as the advantage of the physics-based baseline model are to be expected, as the former originates from the lack of samples, while the latter explicitly applies to all iso-electronic diatomics independent of training set composition.
@@ IV Conclusion
For minimally empirical analytical estimates of energetics among diatomics, one might favor AHA+ours as it provides the correct behavior at short and long distances across the iso-electronic chemical space. It correctly splits in Coulomb and electronic contributions, and it relies only on one calibration point in the entire $(\lambda,d)$-surface.
@@
Future extensions of this research might deal with (i) inclusion of the interatomic distance dependent alchemical force $F_{0}$ into AHA based models, (ii) the extension of chemical space by considering chemical species which are not iso-electronic in total electron count but rather in number of valence electrons, as already examplified for ionic crystals AlchemyAlisa_2016 , covalent bonding Samuel-JCP2016 , and band-gaps of (III)-(V) semi-conductors Samuel2018bandgaps , (iii) extensions to molecules via fragment (amon) based $\Delta$-machine learning Amons ; delta2015 , and (iv) the extension of AHA-baseline $\Delta$-learning to multi-level learning zaspel2018boosting ; heinen2024reducing .
@@ IV.2 Software
Software for the purpose of data generation (e.g. quantum chemistry software) are provided by the Python-packages PySCF PySCF1 ; PySCF2 , basissetexchange bse1 ; bse2 ; bse3 , NumPy numpy and SciPy scipy . Visualizations were created using Matplotlib matplotlib .
@@ IV.3 Data and code availability
The code that produces the figures and findings of this study, in specific the scripts for the generation of DFT data, in addition to the generated data itself, are openly available on Zenodo under zenodo.org/records/13844083.

### v3

# Alchemical harmonic approximation based potential for iso-electronic diatomics: Foundational baseline for $\Delta$-machine learning

@@ Abstract
We introduce the alchemical harmonic approximation (AHA) of the absolute electronic energy for charge-neutral iso-electronic diatomics at fixed interatomic distance $d_{0}$. To account for variations in distance, we combine AHA with this Ansatz for the electronic binding potential, $E(d)=(E_{u}-E_{s})\left(\frac{E_{c}-E_{s}}{E_{u}-E_{s}}\right)^{\sqrt{d/d_{0}}}+E_{s}$, where $E_{u},E_{c},E_{s}$ correspond to the energies of united atom, calibration at $d_{0}$, and sum of infinitely separated atoms, respectively. Our model covers the two-dimensional electronic potential energy surface spanned by distances of 0.7 to 2.5 Åand differences in nuclear charge from which only one single point (with elements of nuclear charge $Z_{1},Z_{2}$ and distance $d_{0}$) is drawn to calibrate $E_{c}$. Using reference data from pbe0/cc-pVDZ, we present numerical evidence for the electronic ground-state of all neutral diatomics with 8, 10, 12, 14 electrons. We assess the validity of our model by comparison to legacy interatomic potentials (Harmonic oscillator, Lennard-Jones, and Morse) within the most relevant range of binding (0.7-2.5 Å), and find comparable accuracy if restricted to single diatomics, and significantly better predictive power when extrapolating to the entire iso-electronic series. We also investigated $\Delta$-learning of the electronic absolute energy using our model as baseline. This baseline model results in a systematic improvement, effectively reducing training data needs for reaching chemical accuracy by up to an order of magnitude from $\sim$1000 to $\sim$100. By contrast, using AHA+Morse as a baseline hardly leads to any improvement, and sometimes even deteriorates the predictive power. Inferring the energy of unseen CO converges to a prediction error of $\sim$0.1 Ha in direct learning, and $\sim$0.04 Ha with our baseline.
@@ II.2 The alchemical harmonic approximation (AHA)
From the constraints in Eqs. 2 and 3, we find for the energy of the AHA:
@@
Knowledge of $F_{c}$ gives a fourth, albeit distance-dependent, constraint on the functional form of $E(\lambda,d)$. With Eq. 5 in mind, we could have picked a different triplet of points or derivatives to determine the parabola, e.g.
@@
However, in addition to being another distance-dependent quantity and consequently requiring the calculation of $\partial E_{s}/\partial\lambda$ (cf. Sec. II.4), both sets of three points rely on the alchemical force $F_{0}(d)$, i.e. any implementation necessitates the calculation of Eq. 19’s derivative w.r.t. $\lambda$ (see below) in addition to $F_{c}$. $F_{c}$ suffers from the errors made in the construction of basis sets which are optimized on energies rather than densities. Thus, for numerical stability, we consider only Eq. 4 because its distance dependence can be expressed without derivatives in $\lambda$.
II.3 The Alchemical Quartic Approximation (AQuA)
Going beyond the quadratic approximation, the next higher polynomial in $E(\lambda)$ must be quartic for symmetry reasons, which we dub the Alchemical Quartic Approximation (AQuA). Using all four previous constraints (Eqs. 2, 3 and 5), we find for the AQuA:
$E(\lambda,d)=\,$ $\left(\frac{F_{0}(d)}{2\lambda_{0}(\lambda_{0}^{2}-\lambda_{u}^{2})}+\frac{E_{u}-E_{0}(d)}{(\lambda_{0}^{2}-\lambda_{u}^{2})^{2}}\right)\lambda^{4}$
$+\left(-2\lambda_{0}^{2}\frac{E_{u}-E_{0}(d)}{(\lambda_{0}^{2}-\lambda_{u}^{2})^{2}}-\frac{F_{0}(d)}{2\lambda_{0}}\frac{\lambda_{0}^{2}+\lambda_{u}^{2}}{\lambda_{0}^{2}-\lambda_{u}^{2}}\right)\lambda^{2}$
$+\left(-\lambda_{u}^{4}\frac{E_{u}-E_{0}(d)}{(\lambda_{0}^{2}-\lambda_{u}^{2})^{2}}+\frac{F_{0}(d)}{2\lambda_{0}}\frac{\lambda_{0}^{2}\lambda_{u}^{2}}{\lambda_{0}^{2}-\lambda_{u}^{2}}\right)$ (14)
However, increasing the polynomial order even further or allowing for fractional powers in $\lambda$ will exacerbate numerical issues and error propagation from constraints, or even remove the possibility to solve for the coefficients analytically. As fifth order polynomials are the highest ones to be analytically solvable, this also excludes more complicated methods like the diatomic energy formula from Ref. michael_hammett_JACS, with powers of $\lambda^{7/3}$ from receiving an analytical treatment.
@@ II.4 Distance-dependence
To model the distance-dependence at some fixed $\lambda_{0}$, we have identified by trial and error the interatomic potential of the diatomic in ground state,
@@
The energies of the neutral atoms $Z_{1},Z_{2},Z_{u}$ are neither $d$-, nor $\lambda$-dependent, and can be precomputed (see Computational Details).
@@
Implicit in this equation is the physical constraint $E_{u}<E_{c}<E_{s}$, i.e. the calibration must never be the united atom which was already a condition for the AHA/AQuA) but also never the infinitely (or very far) separated system! With Eq. 19, only one calibration point $E_{c}$ at $\lambda_{0},d_{0}$ is needed to determine the energy behavior $E(\lambda,d)$ of iso-electronic diatomics for $\lambda<|\lambda_{u}|$ and $d\in[0.7,2.5]$.
@@
Together with the AHA, our interatomic potentials offers a description of energies in diatomics both in $\lambda$ and in distance $d$ from first principles. There are other such approaches e.g. the embedded atom method for crystals of transition metals finnis_1984 ; daw_1984 which employs potentials to model the energy in terms of pair-wise distances, and impurities as deviations in nuclear charge from a so-called "host" electron density, albeit in terms of contributions to the electron density itself. Such an ansatz of jellium plus transmutation is reminiscent of a recent prediction of bonding trends in the context of computational alchemy michael_hammett_JACS . In contrast to (semi-)empirical methods, our interatomic potential in combination with the AHA stems purely from the energy constraints.
@@ III.1 Comparing potentials
Figure 5: Behavior of $E_{s}$ when treated as a fitting parameter evaluated at at different systems $\lambda$ and when precomputed (dashed line). Calibration point is BF ($\lambda_{0}=2$) at $d_{0}=1.2$, $E_{u}$ is precomputed (DFT, pbe0/cc-pVDZ).
Note that conventional potentials, such as Morse, Lennard-Jones or the harmonic potential (harmonic in $d$) all model distance-dependent nuclear repulsion and electronic attraction together—despite the fact that the Coulombic part could have been easily subtracted in order to better focus on the contributions from quantum mechanics (Fig. 3). Especially for short distances, difficulties arise as Coulombic and empirical repulsion do not cancel exactly, e.g. the repulsive term of the Lennard-Jones (LJ) potential ($\propto r^{-12}$) does not describe a physical repulsion. Fitting parameters of the LJ potential minus the Coulomb repulsion to the electronic energy of the diatomic then produces a cancellation of questionable quality as seen in Fig. 3. The advantage of Eq. 19 is its correct and systematic behavior for $d=0$ and $d\rightarrow\infty$, and its dependence on only one calibration point, almost regardless of choice of $d$ and $\lambda$.
Computing parameters via fitting is not necessary for reasonable MAEs of our potential, but naturally, one can consider the quantity $E_{s}$ as fitting parameter to improve Eq. 19’s accuracy. As calibration point, we choose BF ($\lambda_{0}=2$) and $d_{0}=1.2$. The different behavior of a precomputed $E_{s}$ when compared to the fitted one can be found in Fig. 5. However, gaining improved accuracy in the most relevant range of binding $E_{0}(d)$, our potential looses part of its first-principle foundation. In Fig. 4, we present a comparison of our potential with established potentials like the harmonic oscillator (HO), Lennard-Jones (LJ) and the Morse potential, all within the AHA. Again, the calibration point is BF ($\lambda_{0}=2$) at $d_{0}=1.2$. In the fitted potentials (HO, LJ, Morse, our(fit)), there is a clear improvement towards the calibration calculation at $\lambda=2$, but for larger deviations from the symmetric diatomic, the approximation of the AHA worsens (seen in the MAE of ours and ours(fit) which only predict electronic energies). The extreme changes of the other series (HO, LJ, Morse) in MAE stem from a similar source: an improvement of the MAE for unsymmetric charges $Z_{1},Z_{2}$ can be achieved in the fitting routine until the AHA approximation worsens to such a degree that all three potentials can no longer adequately describe the energy shape. Similar behavior can be found throughout the different calibrations (N2, CO, BeNe, LiNa, HeMg, HAl) and the series’ of 8 (Be2, LiB, HeC, HN), 10 (B2, BeC, LiN, HeO, HF) and 12 (C2, BN, BeO, LiF, HeNe, HNa) electrons.
@@ III.2 Machine Learning with AHA as baseline
As a calibration calculation, we again pick BF. Then, the difference between ours and the Morse potential as a baseline model for the prediction of diatomic energies and the reference data at DFT-level of theory (pbe0/cc-pVDZ) can be learned with KRR. We consider DFT to be sufficient for the physical modeling of diatomics while simultaneously retaining low computational cost for the generation of training data. For this, we compute 7 times 1024 points (for the 7 different $\lambda$ and 1024 points for $d\in[0.7,2.5]$   Å), then choose different global and local representations for the diatomics (see Sec. Computational Details). The kernel is Laplacian with the Manhattan norm. To determine hyperparameters, we employ 8-fold cross-validation for different training set sizes $N_{\text{train}}=8,16,32,64\dots,4096$. The size of the test set is $N_{\text{test}}=3072$. We compare these two $\Delta$-learning approaches with direct learning of the CM(n) representation (Fig. 6). Note the different sizes of representations: while CM(n) contains only four numbers per diatomic, cMBDF consists of 40 (times 2 atoms), FCHL19 of 3952 (times 2 atoms) and finally SOAP with 44296 (times 2 atoms)! Especially SOAP scales costly as every element species of the 14-electron diatomics is included, i.e. elements H to Si.
@@ III.3 Out-of-sample prediction of diatomics
We want to determine the MAE of the KRR model when systematically removing one specific diatomic, e.g. CO, from the training set to use as test set. Again, consider the AHA+ours and AHA+Morse baseline models with BF as calibration calculation. Instead of randomly selecting $N_{\text{train}}$ samples, restrict the data in the training set to $\lambda\neq 1$ and in the test set to $\lambda=1$. Since local representations do not work anymore (cf. Eq. 25), we exclusively consider the performance between the CM(n) representations in direct and $\Delta$-learning (inset of Fig. 6). Clearly, all three models approach an average energy prediction early but no accuracy is gained upon further increase of the training set size. However, the baseline model of AHA+ours gives a significant edge over the direct learning or $\Delta$-learning with AHA+Morse. Both the stagnation, as well as the advantage of the physics-based baseline model are to be expected, as the former originates from the lack of samples, while the latter explicitly applies to all iso-electronic diatomics considered here, independent of training set composition.
@@ IV Conclusion
For minimally empirical analytical estimates of energetics among diatomics, one might favor AHA+ours as it provides the correct behavior at short and long distances across the iso-electronic chemical space. It correctly splits in Coulomb and electronic contributions, and it relies only on one calibration point in the $(\lambda,d)$-surface.
@@
Future extensions of this research might deal with (i) inclusion of the interatomic distance dependent alchemical force $F_{0}$ into AHA based models, (ii) the extension of chemical space by considering chemical species which are not iso-electronic in total electron count but rather in number of valence electrons, as already exemplified for ionic crystals AlchemyAlisa_2016 , covalent bonding Samuel-JCP2016 , and band-gaps of (III)-(V) semi-conductors Samuel2018bandgaps , (iii) extensions to molecules via fragment-based $\Delta$-machine learning delta2015 like e.g. atoms-in-molecules (amons) Amons , (iv) the extension of AHA-baseline $\Delta$-learning to multi-level learning zaspel2018boosting ; heinen2024reducing , and (v) a generalization of our interatomic potential beyond the ground state.
@@ IV.2 Software
Software for the purpose of data generation (e.g. quantum chemistry software) are provided by the Python-packages PySCF PySCF1 ; PySCF2 and specifically its internal function dft.ks, basissetexchange bse1 ; bse2 ; bse3 , NumPy numpy and SciPy scipy . Visualizations were created using Matplotlib matplotlib .
@@ IV.3 Data and code availability
The code that produces the figures and findings of this study, specifically the scripts for generation of DFT data, in addition to the data itself, are openly available on Zenodo under www.zenodo.org/records/13844083.
Below we give the energies of elements H to Si employed for the generation of data in AHA+ours. Their spin has been kept either zero or one, depending on even or odd number of electrons, respectively. A minimal working example for $Z=3$ with Kohn-Sham-DFT, basis set cc-pVDZ ccpvnz1 ; ccpvnz2 ; ccpvnz3 and the pbe0-functional PBE0 ; PBE01 ; PBE02 using PySCF PySCF1 ; PySCF2 reads:
from pyscf import gto, dft
import basis_set_exchange as bse
Z = 3
basis_set = ’cc-pVDZ’
xc = ’pbe0’
basis = bse.get_basis(basis_set, fmt=’nwchem’)
mol = gto.M(atom = str(Z)+’ 0 0 0’,
            charge=0,
            spin=Z%2,
            basis = basis)
mf = dft.KS(mol)
mf.xc = xc
mf.kernel()
This results in the numerical values in Tab. 1.
Nuclear charge $Z$ Energy [Ha]
0 0.0000
1 -0.5002
2 -2.8876
3 -7.4661
4 -14.6351
5 -24.6159
6 -37.7241
7 -54.4864
8 -74.8817
9 -99.6356
10 -128.8050
11 -162.1752
12 -199.9610
13 -242.2462
14 -289.1975
Table 1: Energies of elements H to Si employed for the generation of data with AHA+ours